OpenStudy (anonymous):

Let f(x) = x – 1 and g(x) = 2x + 3. What is f(g(x)) ? 2x + 2 2x + 1 3x + 2 x + 4

OpenStudy (anonymous):

\[f(g(x)) = (2x + 3) - 1 = 2x + 3 - 1 = 2x +2\]

OpenStudy (anonymous):

Put x in f(x) with 2x + 3: \[f(2x+3) = (2x +3) -1 = 2x + 2\]

OpenStudy (anonymous):

\(f(2x+3) = f(g(x))\)..